Could you please help me?
Find an equation of the line having the given slope and containing the given point. m = 2/3, (8,-4)
The equation of the line y= ? I have tried and tried and still not sure what I am doing. Thanks for all your help!

Answer by Alan3354(69443)

Use y = mx + b and the point to find b, the y-intercept.
---------
y = -4, x = 8
-4 = (2/3)*8 + b
b = 4/3
--------
y = (2/3)x + 4/3 in slope-intercept form
2x - 3y = -4 in standard form